Make New Entry,
Make Followup Entry
User name J. Wexler
Log entry time 08:00:02 on August26,2009
Entry number 286095
keyword=Shift summary 8/26 owl
Shift summary for Owl shift, 8/26:
12AM - No beam, currently tuning.
12:35 - Change to thin tantalum
12:40 - Changed magnets to the PREX tune.
12:50 - 5 uA beam on the flat, 1cm PREX detector (runs 4331-4).
2:00 - Changed to thick tantalum target.
2:05 - Request MCC ramp up the beam on the thick tantalum target to 100uA. Begin the MCCs program for current ramping
2:57 - Switch to empty target, per MCC request
3:26 - Switch to thick tantalum target, per MCC request
3:28 - Switch to empty target, per MCC request
3:30 - MCC ramps current up to 100 uA
4:20 - Reach 100 uA, switch in thick tantalum.
5:30 - MCC informs us that achieving 100uA on the thick target will require changing ion chamber trip levels. We agree to recieve 30 uA.
5:35 - Recieve 30uA for oversampling tests on the lumis (4335-8)
6:13 - Stop lumi tests, call for access to change detector configurations
6:55 - Switch to 45 deg. angled PREX detector configuration. MCC says beam is currently down.
7:30 - Recieve beam from the MCC, test the 45 deg. config. (4339-4341).